a sofa is on sale for 64 % off
the sale price is $316.80
what is the regular price

Answer by jim_thompson5910(35256)

Let x = regular price

The equation is then

x - 0.64x = 316.8

Now solve for x

x - 0.64x = 316.8


0.36x = 316.8


x = 316.8/0.36


x = 880


So the regular price is $880